# Rigidity for Patterson–Sullivan systems with applications to random walks and entropy rigidity

### v2

@@ 10.1. Contracting isometries
(CTG) $\Gamma<\operatorname{\mathsf{Isom}}(X)\text{ discrete with a contracting isometry.}$.
@@ 10.5. Verification of PS-system
We further show that the PS-system in Theorem 10.11 is well-behaved under some condition related to a saturation of $\partial_{H}X$; w call $c\in\partial_{H}X$ saturated if for any $c^{\prime}\in\overline{X}\smallsetminus\{c\}$, $\|c-c^{\prime}\|_{\infty}=+\infty$.
@@
[…]
@@ 10.7. Invisible locus
Fix a metric on $\overline{X}$ which generates the topology. Property (PS3) is implied by Property (PS5). To see Property (PS5), let $\{\gamma_{n}\}\subset\Gamma$ and $R_{n}\to+\infty$ be sequences such that $[\partial_{H}X\smallsetminus\gamma_{n}^{-1}\operatorname{\mathcal{O}}_{R_{n}}(\gamma_{n})]\to Z$ with respect to the Hausdorff distance. After passing to a subsequence, we may assume that $\gamma_{n}^{-1}o\to z\in\partial_{H}X$. By Proposition 10.15, for any $h_{1},\ldots,h_{m}\in\Gamma$, there exists an $\alpha$-contracting isometry $g\in\Gamma$ such that $z,h_{1}z,\ldots,h_{m}z\notin\partial^{+}(\langle g\rangle o)$, for some $\alpha\geq 0$. Then Property (PS5) is a conseqeunce of Lemma 10.16 and Lemma 10.17. ∎
@@ 10.9. Properties of contracting tails
Property (PS8) says that shadows converging to a generic point have diameter decaying to 00. This can be observed from contracting tails. Recall that $c\in\partial_{H}X$ is saturated if for any $c^{\prime}\in\overline{X}\smallsetminus\{c\}$, $\|c-c^{\prime}\|_{\infty}=+\infty$.
@@ 12.2. Random walks on mapping class groups and Teichmüller spaces
Let $d_{w}$ be a word metric on $\Gamma$ with respect to a finite generating set. Fix $\epsilon>0$. It is easy to see that there exists $R_{0}>0$ such that for any $\gamma\in\Gamma$ and $x\in\gamma\left(\partial_{\infty}\Gamma\smallsetminus\overline{B_{\epsilon}(\gamma^{-1})}\right)$, any geodesic ray $[\operatorname{id},x)\subset\Gamma$ with respect to $d_{w}$ intersects the $\operatorname{d}_{w}$-ball of radius $R_{0}$ centered at $\gamma$.
Let $\gamma\in\Gamma$ and $x\in\gamma\left(\partial_{\infty}\Gamma\smallsetminus\overline{B_{\epsilon}(\gamma^{-1})}\right)$. Fix a geodesic ray $[\operatorname{id},x)\subset\Gamma$ and for each $n\geq 1$, let $\gamma_{n}\in[\operatorname{id},x)$ be such that $\operatorname{d}_{w}(\operatorname{id},\gamma_{n})=n$. By [Miy13],
@@ 12.3. Random walks on discrete subgroups of Lie groups
[…]
@@
If $\mu$ is a coarse $\phi$-PS measure for $\Gamma$ on $\operatorname{\mathcal{F}}$ of dimension $\delta$ and the measures $\nu$, $\mu$ are not singular, then:
@@
If $\Gamma$ is word hyperbolic (as an abstract group) and contains a unipotent element of $\operatorname{\mathsf{G}}$, then $\nu$ is singular with respect to every coarse Iwasawa PS-measure on $\operatorname{\mathcal{F}}$.

### v3

@@ 10.1. Contracting isometries
(CTG) $\Gamma<\operatorname{\mathsf{Isom}}(X)\text{ discrete with a contracting isometry.}$
@@ 10.5. Verification of PS-system
[…]
@@
$\operatorname{\mathcal{C}}(\alpha,L):=\{\gamma\in\Gamma:(o,\gamma o)\text{ has an }(\alpha,L)\text{-contracting tail}\}.$
@@ 10.7. Invisible locus
Fix a metric on $\overline{X}$ which generates the topology. Property (PS3) is implied by Property (PS5). To see Property (PS5), let $\{\gamma_{n}\}\subset\Gamma$ and $R_{n}\to+\infty$ be sequences such that $\partial_{H}X\smallsetminus\gamma_{n}^{-1}\operatorname{\mathcal{O}}_{R_{n}}(\gamma_{n})]\to Z$ with respect to the Hausdorff distance. After passing to a subsequence, we may assume that $\gamma_{n}^{-1}o\to z\in\partial_{H}X$. By Proposition 10.15, for any $h_{1},\ldots,h_{m}\in\Gamma$, there exists an $\alpha$-contracting isometry $g\in\Gamma$ such that $z,h_{1}z,\ldots,h_{m}z\notin\partial^{+}(\langle g\rangle o)$, for some $\alpha\geq 0$. Then Property (PS5) is a conseqeunce of Lemma 10.16 and Lemma 10.17. ∎
@@ 10.9. Properties of contracting tails
[…]
@@ 12.2. Random walks on mapping class groups and Teichmüller spaces
Let $d_{w}$ be a word metric on $\Gamma$ with respect to a finite generating set. Fix $\epsilon>0$. It is easy to see that there exists $R_{0}>0$ such that for any $\gamma\in\Gamma$ and $x\in\gamma\left(\partial_{\infty}\Gamma\smallsetminus\overline{B_{\epsilon}(\gamma^{-1})}\right)$, any geodesic ray $\operatorname{id},x)\subset\Gamma$ with respect to $d_{w}$ intersects the $\operatorname{d}_{w}$-ball of radius $R_{0}$ centered at $\gamma$.
Let $\gamma\in\Gamma$ and $x\in\gamma\left(\partial_{\infty}\Gamma\smallsetminus\overline{B_{\epsilon}(\gamma^{-1})}\right)$. Fix a geodesic ray $\operatorname{id},x)\subset\Gamma$ and for each $n\geq 1$, let $\gamma_{n}\in[\operatorname{id},x)$ be such that $\operatorname{d}_{w}(\operatorname{id},\gamma_{n})=n$. By [Miy13],
@@ 12.3. Random walks on discrete subgroups of Lie groups
As a higher rank analogue of critical exponent, Quint introduced the notion of growth indicator on $\Gamma$ [Qui02b]. Fixing any norm $\left\|\cdot\right\|$ on $\mathfrak{a}$, the growth indicator of $\Gamma$ is the function $\psi_{\Gamma}:\mathfrak{a}\to\operatorname{\mathbb{R}}\cup\{-\infty\}$ defined as follows: for $u\neq 0$,
@@
[…]
@@
[…]